# Neural Green’s Operators for Parametric Partial Differential Equations

### v4

@@ 4.3.6 NGO-based Preconditioners
    Next, we compute an SVD of the inner matrix: $\mathbf{L}^{\top}\mathbf{A}[\theta]\mathbf{L}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, with $\mathbf{U}$ and $\mathbf{V}$ orthogonal matrices and $\bm{\Sigma}=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{n})$:
    $\hat{G}[\theta](\mathbf{x},\mathbf{x}^{\prime})$ $\approx\sum_{m,n}\tilde{\phi}_{m}(\mathbf{x})\left(\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}\right)_{mn}\tilde{\phi}_{n}(\mathbf{x}^{\prime}).$ (4.19)
    We can then define two additional orthonormal bases, by merging the factors $\mathbf{U}$ and $\mathbf{V}$ into the basis $\tilde{\phi}_{m}$:
    $\tilde{\phi}^{(U)}_{n}$ $=\sum_{m}\tilde{\phi}_{m}\mathbf{U}_{mn},$ (4.20a)
    $\tilde{\phi}^{(V)}_{n}$ $=\sum_{m}\tilde{\phi}_{m}\mathbf{V}_{mn},$ (4.20b)
    $\hat{G}[\theta](\mathbf{x},\mathbf{x}^{\prime})$ $\approx\sum_{m}\tilde{\phi}^{(U)}_{m}(\mathbf{x})\sigma_{m}\tilde{\phi}^{(V)}_{m}(\mathbf{x}^{\prime}).$ (4.20c)
Now, we focus on the orthonormal basis functions $\tilde{\phi}^{(U)}_{i}$: these are the left singular functions corresponding to the largest singular values of $\hat{G}$, and will therefore correspond to the smallest singular values of the differential operator $\mathbf{C}$. We now want a preconditioner that, when given a vector $\mathbf{u}$ corresponding to a function $u(\mathbf{x})$, amplifies the low-frequency components in $u$ that correspond to the basis functions $\tilde{\phi}^{(U)}$, while preserving the higher frequency components. For a function $u$, its decomposition into low and high frequencies is as follows:
$u(\mathbf{x})$ $=\underbrace{\sum_{i}\left\langle\tilde{\phi}^{(U)}_{i},u\right\rangle\tilde{\phi}^{(U)}_{i}(\mathbf{x})}_{\text{low frequencies}}+\underbrace{\left[u(\mathbf{x})-\sum_{i}\left\langle\tilde{\phi}^{(U)}_{i},u\right\rangle\tilde{\phi}^{(U)}_{i}(\mathbf{x})\right]}_{\text{high frequencies}},$ (4.21)
where the orthonormality of the basis $\tilde{\phi}^{(U)}$ is crucial in making the separation easy to compute. To do this separation in the discrete setting, we sample the functions $\tilde{\phi}^{(U)}_{i}$, producing a matrix $\bm{\Phi}\in\mathbb{R}^{N\times n}$, and replace the $L^{2}$ inner products in $\Omega$ by a quadrature rule with weights $\mathbf{W}\in\mathbb{R}^{N}$:
$\mathbf{u}$ $=\underbrace{\bm{\Phi}\bm{\Phi}^{\top}\text{diag}(W)\mathbf{u}}_{\text{low frequencies}}+\underbrace{\left[\mathbf{u}-\bm{\Phi}\bm{\Phi}^{\top}\text{diag}(W)\mathbf{u}\right]}_{\text{high frequencies}}.$ (4.22)
Now, the preconditioner can amplify the low frequencies as $r\bm{\Phi}\mathbf{\Sigma}\bm{\Phi}^{\top}\text{diag}(W)\mathbf{u}$: this amplifies each frequency proportionally to how strongly $\mathbf{C}$ dampens that frequency, where the factor $r$ remains to be chosen. If we add the high frequency components to this, we obtain a preconditioner:
$\mathbf{M}^{-1}\mathbf{u}$ $=\bm{\Phi}\mathbf{\Sigma}\bm{\Phi}^{\top}\text{diag}(W)\mathbf{u}+r\left[\mathbf{u}-\bm{\Phi}\bm{\Phi}^{\top}\text{diag}(W)\mathbf{u}\right]=r\mathbf{u}+\bm{\Phi}\left(\mathbf{\Sigma}-r\mathcal{I}\right)\bm{\Phi}^{\top}\text{diag}(W)\mathbf{u}.$ (4.23)
In practice, it may not be the case that the NGO learns the first $n=144$ modes accurately. Therefore, instead of taking all basis functions, we only take the first $k<n$ basis functions, resulting in the preconditioner:
$\mathbf{M}_{k,r}^{-1}\mathbf{u}$ $=r\mathbf{u}+\bm{\Phi}_{:,1:k}\left(\mathbf{\Sigma}_{1:k}-r\mathcal{I}\right)\bm{\Phi}_{:,1:k}^{\top}\text{diag}(W)\mathbf{u}.$ (4.24)
The value of $r$ should now be such that the combined effect of the preconditioner and the differential operator, i.e. the matrix $\mathbf{CM}^{-1}$, amplifies low frequencies and high frequencies approximately equally. One reasonable choice therefore is $r\propto\sigma_{k+1}$, which is the lowest amplification factor among the high frequencies. In the experiments in this section, we chose $r=\frac{\sigma_{k+1}}{2}$.
Note that the bulk of this process is independent of the resolution of the ‘full-order’ numerical scheme. Specifically, for an NGO with $n$ basis functions preconditioning an $N\times N$ linear system, we can see the following:
  • The orthogonalization of the basis requires only a Cholesky decomposition of an $n\times n$ matrix. In fact, if the basis $\phi_{m}$ has tensor structure (as is the case here), this can be factorized further. This step depends only on the basis, and is therefore independent of the parameters $\theta$, meaning the decomposition only needs to be done once.
  • The second expensive step is the SVD, which is performed on the matrix $\mathbf{L}^{-1}\mathbf{AL}^{-\top}$, which is also $n\times n$.
  • The preconditioner requires evaluating $\bm{\Phi}^{(U)}$, which requires evaluating the basis functions on the NN points (producing an $N\times n$ matrix) and multiplying by the $n\times n$ matrix $\mathbf{L}^{-\top}\mathbf{U}$.
  • Finally, the resulting preconditioner, although it is $N\times N$ and dense, does not need to be stored in dense form since its action on a vector $v$ can be computed easily:
    $\mathbf{M}^{-1}\mathbf{u}$ $=r\mathbf{u}+\underbrace{\bm{\Phi}^{(U)}}_{N\times k}\underbrace{\left[\left(\mathbf{\Sigma}-r\mathcal{I}\right)\left(\bm{\Phi}^{(U)}\right)^{\top}\text{diag}(W)\right]}_{k\times N}\mathbf{u},$ (4.25)
    which means that after precomputing the two matrices indicated, $\mathbf{M}^{-1}\mathbf{u}$ can be computed in $\mathcal{O}(kN)$ time.
Since the NGO uses a certain function basis, the preconditioner only affects the components of the solution that lie in this basis, which are here associated with the low frequencies of the solution. This means that additional preconditioning can be introduced to accelerate convergence of the high-frequency components. One way this can be achieved is using multiplicative preconditioning, with a block-Jacobi preconditioner as a pre- and post-smoother. In this case, the combined preconditioner is expressed most naturally not as a matrix, but as a linear function $\mathbf{r}\mapsto\mathbf{u}^{(3)}$ comprised of the following three steps:
$\mathbf{u}^{(1)}$ $=\mathbf{M}_{\text{block-Jacobi}}^{-1}\mathbf{r},$ (4.26a)
$\mathbf{u}^{(2)}$ $=\mathbf{u}^{(1)}+\mathbf{M}_{\text{NGO}}^{-1}\left(\mathbf{r}-\mathbf{A}\mathbf{u}^{(1)}\right),$ (4.26b)
$\mathbf{u}^{(3)}$ $=\mathbf{u}^{(2)}+\mathbf{M}_{\text{block-Jacobi}}^{-1}\left(\mathbf{r}-\mathbf{A}\mathbf{u}^{(2)}\right).$ (4.26c)
Figure 12 shows the convergence behavior of GMRES. Here, the preconditioners were constructed from the data NGO trained in Section 4.3.1. For each of the preconditioner choices, 1000 test problems were discretized with finite differences and then solved with GMRES with a tolerance of $10^{-5}$, with GMRES restarting every $20$ iterations. This plot shows that using an NGO preconditioner greatly accelerates convergence compared to the un-preconditioned case, and also results in faster convergence than when using only a block-Jacobi preconditioner. However, the fastest convergence is obtained when both preconditioners are used in the way described above. Table 4 shows the mean number of GMRES iterations for different combinations of preconditioners, and the corresponding improvement over un-preconditioned GMRES. Note that the addition of an NGO preconditioner speeds up the convergence of GMRES by more than a factor 20, also when comparing the block-Jacobi preconditioner to the combined preconditioner.
Figure 12: Convergence behavior of GMRES for $100\times 100$ finite-difference discretizations of the diffusion equation, in four different cases: without any preconditioner, with block-Jacobi preconditioner, with NGO preconditioner, and with a combined preconditioner.
Table 4: Mean number of GMRES iterations required for convergence with different preconditioners.
Preconditioner GMRES iterations Improvement factor
None 1298 -
Block-Jacobi 599 2.2
NGO 62 20.9
NGO + block-Jacobi 29 44.5
@@ 5.1 Time-Dependent Diffusion
where $\Omega=[0,L]^{2}$ and $\mathcal{T}=[0,T]$, which is a direct time-dependent generalization of the steady diffusion problem as treated in Section 4. The differences with respect to the static problem are that the operator in (5.1) contains a time derivative $\partial/\partial t$, the input functions $\theta$, $f$, $\eta$ and $g$ are functions of space and time $(\mathbf{x},t)$, and there is an initial condition $u=u_{0}$. Instead of resolving the Green’s operator corresponding to (5.1) on the full time domain $\mathcal{T}$, we restrict ourselves to a time slab of length $\Delta t$, as shown in Figure 13.
Figure 13: The space-time domain $\Omega\times\mathcal{T}$, with Neumann boundaries $\Gamma_{\mathrm{N}}$ on the top and bottom, Dirichlet boundaries $\Gamma_{\mathrm{D}}$ on the sides, and the initial condition boundary $t=0$ on the left side.
@@ 5.1.2 Inductive Bias: Time Stepping Stability
Figure 14: Effect of adding the Nitsche-stabilized Neumann series preconditioning to a model NGO, and the norm scaling layer (5.20) to a data NGO, on their (a) mass matrix $\mathbf{M}_{\mathrm{rr}}$-generalized eigenvalue spectrum and (b) relative time step $L^{2}$ test error over time. Projection and stabilized FEM are included as reference in black and gray, respectively. In (a), points correspond to the spectrum of a single matrix corresponding to a manufactured $\theta$ (manufactured as in dataset D). In (b), lines and shaded areas are, respectively, 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=\tau/T=1$.
@@
Figure 15: Effect of adding the conservation correction (5.26) to a data NGO on the errors of its predicted (a) $m(t)=\int_{\Omega}u(\mathbf{x},t)d\mathbf{x}$ and (b) energy $E(t)=1/2\int_{\Omega}u^{2}(\mathbf{x},t)d\mathbf{x}$ over time. Projection and stabilized FEM are included as reference in black and gray, respectively. All models are trained on dataset D (Appendix B). Lines and shaded areas are, respectively, averages and 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=\tau/T=1$.
@@
Figure 16: Effect of adding the conservation correction (5.26) to a data NGO. Projection and stabilized FEM are included as reference in black and gray, respectively. All models are trained on dataset D (Appendix B). Lines and shaded areas are, respectively, averages and 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=\tau/T=1$.
@@ 5.1.3 Results
In this section, we perform the same test as considered in Figure 8, except that here, we compare NGOs to other neural operators in their generalization across test data length as well as time scales. Additionally, we compare the long term stability of NGOs against the same models from literature. The architectural details of the used models are essentially identical to those corresponding to Figure 8, except that the models now also use quadrature and basis functions in space-time. Specifically, the DeepONet, VarMiON, data NGO, data-free NGO and model NGO now use a 2 linear $\times$ 10 cubic $\times$ 10 cubic B-spline basis in space-time $(t,x,y)$ and all models use a quadrature grid of approximately $2\times 100\times 100$ points. The FNO uses $2\times 10\times 10$ Fourier modes, and the standalone NN (U-Net) uses 3-dimensional convolutions. The VarMiON is now limited to using only $2\times 5\times 5$ uniform quadrature points, because of the limited parameter budget of $3\cdot 10^{4}$ trainable parameters. Details of the model architectures are highlighted in Appendix C.2. The models have been trained on dataset D, described in Appendix B.2, using the training procedure described in Appendix D.2.
@@
Figure 17: Single time step relative $L^{2}$ test error versus (a) time scale $\tau/T$ and (b) length scale $\lambda/L$ for a bare NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO. All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO) use the same 2 linear $\times$ 10 cubic $\times$ 10 cubic B-spline basis in space-time $(t,x,y)$. Points and error bars are, respectively, averages and 95% confidence intervals on 1000 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=1$ (in (a)) and with $\tau/T=1$ (in (b)).
@@
Figure 18: Relative time step $L^{2}$ test error over time on (a) in distribution data with time and length scale $\tau/T=\lambda/L=1$ and (b) out of distribution data with $\tau/T=\lambda/L=0.1$, for a bare NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO (an FNO is also added for comparison). All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO) use the same 2 linear $\times$ 10 cubic $\times$ 10 cubic B-spline basis, and are thus limited by the same $L^{2}$ projection error lower bound, indicated in black. Points and shaded areas are, respectively, averages and 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\tau/T=\lambda/L=1$ (in (a)) and with $\tau/T=\lambda/L=0.1$ (in (b)).

### v5

@@ 4.3.6 NGO-based Preconditioners
    Expanding the resulting coefficients in terms of the basis $\phi_{m}$, denoted by the operator $\mathbf{P}_{ij}$, represents a prolongation operator.
In this way, the application of preconditioners inferred from NGOs can be applied algebraically to any Krylov subspace method. Furthermore, the multiplicative combination with a block-Jacobi preconditioner works as follows: after the application of the NGO-based preconditioner, the block-Jacobi preconditioner is applied as a correction:
$\tilde{\mathbf{x}}$ $=\mathbf{P}_{\text{NGO}}(\mathbf{v})$ (4.22a)
$\tilde{\mathbf{r}}$ $=\mathbf{v}-\mathbf{C}\tilde{\mathbf{x}}$ (4.22b)
[…]
Preconditioning with DeepONets
To use a DeepONet as a preconditioner, we follow the same procedure as described in Kopaničáková et al. [kopaničáková2025deeponet], although our setup is slightly different since we apply the preconditioner to a finite-difference discretization, rather than a finite-element discretization. This means that in order to compute the right-hand side vector for the DeepONet, which consists of the source terms evaluated on the sensor nodes of the DeepONet, $f$ is interpolated from the finite-difference discretization points onto the sensor nodes using nearest-neighbor interpolation (i.e. the inputs to the DeepONet are taken to be the pointwise values on the discretization points $\mathbf{x}_{i}$ closest to each of the $12\times 12$ sensor nodes of the DeepONet). While this step technically introduces an error in the interpolation, we believe this error to be negligible overall as the mesh size of $h=\frac{1}{99}$ means that the difference between the DeepONet’s sensor nodes and the actually sampled locations is small compared to the length scales of the input.
There are two more important details in the application of DeepONets as preconditioners:
  • The first is that while DeepONets are passed the individual source terms separately, in a discretization these terms are all combined into one right-hand-side vector. Since most of the entries in the right-hand-side vector follow from the volumetric source term $f$ and not the boundary source term $\eta$ (note that the Dirichlet boundary data $g$ is always zero in the data set used here), the choice is made here to always set $\eta\equiv 0$ and treat the right-hand-side vector $\mathbf{v}$ as if it is entirely the result of the volumetric source term $f$. While this may also introduce an error, there is no clear way to “separate” a right-hand side vector into contributions from the volumetric and boundary source terms and this limitation is not addressed in the original work.
  • Secondly, since DeepONets are nonlinear functions, they may produce unexpected outputs when the right-hand-side vector $\mathbf{v}$ is of a different order of magnitude than what is present in the training data. This is avoided by scaling the input vector so that its norm equals the average $L^{2}$-norm of the vectors $\mathbf{f}$ in the training data, written here as $r_{0}$. This enforces that the DeepONet-based preconditioner, while not linear, is positively homogeneous.
Algorithmically, applying the DeepONet-based preconditioner to a vector $\mathbf{v}\in\mathbb{R}^{9800}$ can be expressed in pseudocode as follows:
$\mathbf{f}$ $=\text{Interpolate}(\mathbf{v})$ nearest-neighbor interpolation (4.23a)
$r$ $=\frac{\|\mathbf{f}\|}{r_{0}}$ normalization (4.23b)
$\mathbf{c}$ $=\text{BranchNet}(\bm{\theta},\tfrac{1}{r}\mathbf{f},\bm{\eta}\equiv 0)$ apply branch net (4.23c)
$\hat{\mathbf{u}}_{i}$ $=\sum_{m}\mathbf{c}_{m}\phi_{m}(\mathbf{x}_{i})$ evaluate solution on grid points (4.23d)
$\mathbf{C}^{-1}\mathbf{v}$ $\approx\mathbf{P}_{\text{DeepONet}}(\mathbf{v}):=r\,\hat{\mathbf{u}}$. undo normalization (4.23e)
The multiplicative combination with the block-Jacobi preconditioner is done in the same way as for NGOs, i.e. as in (4.22).
Results
To test the efficacy of NGOs and DeepONets as preconditioners, we consider 1000 realizations of the diffusion problem. We solve the discretized equations with F-GMRES [52] restarted every 50 inner iterations, as well as unrestarted F-GMRES (written here as F-GMRES(50) and F-GMRES($\infty$), respectively). However, F-GMRES(50) was found to diverge when used with a DeepONet preconditioner, and therefore the DeepONet preconditioning results are only available for unrestarted F-GMRES. Since the NGO preconditioner can also be applied to other linear solvers, the equations were also solved using the regular GMRES algorithm [51] (both restarted and unrestarted), as well as Bi-CGSTAB [57]. The average number of iterations needed for convergence is shown in Table 4. The linear systems are solved with a relative tolerance of $10^{-8}$. Right preconditioning is used in all cases.
Table 4: The average number of iterations that various iterative linear solvers needed to solve discretized diffusion equations. These averages are computed over 1000 realizations of diffusion problems.
Solver Preconditioner
None Blk. Jac. Blk. Jac. + NGO Blk. Jac. + DeepONet
F-GMRES($\infty$) 482.9 293.7 50.0 274.9
F-GMRES(50) 1616.2 719.4 50.0 Fail
GMRES($\infty$) 482.9 293.7 50.0 -
GMRES(50) 1616.2 719.4 50.0 -
Bi-CGSTAB 360.1 213.2 29.4 -
The convergence behavior of different solvers with various preconditioners is shown in Figure 12. This figure shows that the NGO-based and DeepONet-based preconditioners accelerate F-GMRES in different ways: while the NGO-based preconditioner results in consistently fast convergence, the DeepONet causes F-GMRES to converge slowly at first but converge more quickly in later iterations.
Figure 12: Convergence behavior of F-GMRES($\infty$) and Bi-CGSTAB with various preconditioners.
Note that in Table 4, some combinations of solvers and preconditioners require the exact same number of iterations on average. This is due to two reasons:
  • Firstly, F-GMRES is equivalent to regular GMRES when used with a linear preconditioner, or with no preconditioner at all. Therefore, the convergence behavior of F-GMRES($\infty$) and F-GMRES(50) are equivalent to that of GMRES($\infty$) and GMRES(50), respectively (with the exception of the DeepONet-based preconditioner).
  • Secondly, with the block-Jacobi + NGO preconditioner, F-GMRES and GMRES very consistently converge in approximately 50 iterations, meaning that the restarted variants converge without ever restarting and therefore behave identically to their unrestarted counterparts. In particular, the “50.0” entries in Table 4 are not erroneous and are confirmed by Figure 12 which shows unrestarted F-GMRES converge in 50 iterations with the NGO-based preconditioner.
An important observation is that F-GMRES($\infty$) with the block-Jacobi + DeepONet preconditioner needs almost as many iterations on average as F-GMRES($\infty$) with only a block-Jacobi preconditioner. In other words, the inclusion of the DeepONet is not very productive. We suspect that this is due to the worse generalizability of DeepONets compared to NGOs: as the F-GMRES algorithm iterates, it may apply the preconditioner to vectors that don’t correspond directly to right-hand side terms in the training data, i.e. that lie out of the training distribution. As shown in Table 2, DeepONets are found not to generalize well beyond the training data set, and this may explain their poor performance as preconditioners in this test case. NGOs, on the other hand, were found to generalize relatively well, and indeed they are also highly effective preconditioners.
We must note that if generalization is indeed the cause of this, then the poor preconditioner performance of DeepONets could be avoided by training on a richer data set, so that the preconditioner performance relies less on the generalizability of the model. As such, the results found here do not contradict those of Kopaničáková et al. [kopaničáková2025deeponet], who found DeepONets to make for effective preconditioners. Nevertheless, our results show that NGOs can learn effective preconditioners from the same training data. Furthermore, their linearity allows NGOs to precondition linear solvers other than F-GMRES. This is relevant as Table 4 shows Bi-CGSTAB to be a more efficient algorithm for the diffusion problem considered here, and Bi-CGSTAB cannot be used with a nonlinear preconditioner such as one derived from a DeepONet.
Remark 7.
In their work, Kopaničáková et al. [kopaničáková2025deeponet] also consider a second type of machine learning-based preconditioner, called the trunk-basis (TB) approach. There, the learned basis of the DeepONet is used to construct the restriction and prolongation operators $\mathbf{R}$ and $\mathbf{P}$, after which the preconditioner is constructed without using the branch network of the DeepONet. In this work, the DeepONet and NGO models both utilize the same basis, thereby making the trunk-basis approach trivially equivalent between both models. As such, the trunk-basis approach is omitted from the comparison.
Remark 8.
Our construction of NGO-based preconditioners can be viewed as a coarse-level preconditioner within a classical two-level framework, in line with the approach of Nikolopoulos et al. [41] and, in the context of model order reduction, with the work of Pasetto et al. [42].
@@ 5.1 Time-Dependent Diffusion
where $\Omega=[0,1]^{2}$ and $\mathcal{T}=[0,T]$, which is a direct time-dependent generalization of the steady diffusion problem as treated in Section 4. The differences with respect to the static problem are that the operator in (5.1) contains a time derivative $\partial/\partial t$, the input functions $\theta$, $f$, $\eta$ and $g$ are functions of space and time $(\mathbf{x},t)$, and there is an initial condition $u=u_{0}$. Instead of resolving the Green’s operator corresponding to (5.1) on the full time domain $\mathcal{T}$, we restrict ourselves to a time slab of length $\Delta t$, as shown in Figure 13.
Figure 13: The space-time domain $\Omega\times\mathcal{T}$, with Neumann boundaries $\Gamma_{\mathrm{N}}$ on the top and bottom, Dirichlet boundaries $\Gamma_{\mathrm{D}}$ on the sides, and the initial condition boundary $t=0$ on the left side. We used $L=T=1$.
@@ 5.1.2 Inductive Bias: Time Stepping Stability
Figure 14: Effect of adding the Nitsche-stabilized Neumann series preconditioning to a model NGO, and the norm scaling layer (5.20) to a data NGO, on their (a) mass matrix $\mathbf{M}_{\mathrm{rr}}$-generalized eigenvalue spectrum and (b) relative time step $L^{2}$ test error over time. Projection and stabilized FEM are included as reference in black and gray, respectively. In (a), points correspond to the spectrum of a single matrix corresponding to a manufactured $\theta$ (manufactured as in dataset D). In (b), lines and shaded areas are, respectively, 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=\tau/T=1$ (where we used $L=T=1$).
@@
Figure 15: Effect of adding the conservation correction (5.26) to a data NGO on the errors of its predicted (a) $m(t)=\int_{\Omega}u(\mathbf{x},t)d\mathbf{x}$ and (b) energy $E(t)=1/2\int_{\Omega}u^{2}(\mathbf{x},t)d\mathbf{x}$ over time. Projection and stabilized FEM are included as reference in black and gray, respectively. We note that the apparent mass conservation error in the projected solution arises from quadrature inaccuracies and would otherwise vanish. All models are trained on dataset D (Appendix B). Lines and shaded areas are, respectively, averages and 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=\tau/T=1$ (where we used $L=T=1$).
@@
Figure 16: Effect of adding the conservation correction (5.26) to a data NGO. Projection and stabilized FEM are included as reference in black and gray, respectively. All models are trained on dataset D (Appendix B). Lines and shaded areas are, respectively, averages and 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=\tau/T=1$ (where we used $L=T=1$).
@@ 5.1.3 Results
In this section, we perform the same test as considered in Figure 8, except that here, we compare NGOs to other neural operators in their generalization across test data length as well as time scales. Additionally, we compare the long term stability of NGOs against the same models from literature. The architectural details of the used models are essentially identical to those corresponding to Figure 8, except that the models now also use quadrature and basis functions in space-time. Specifically, the DeepONet, VarMiON, data NGO, data-free NGO and model NGO now use a 2 linear $\times$ 10 cubic $\times$ 10 cubic B-spline basis in space-time $(t,x,y)$ and all models use a quadrature grid of approximately $2\times 100\times 100$ points. The FNO uses $2\times 10\times 10$ Fourier modes, and the standalone NN (U-Net) uses 3-dimensional convolutions. The VarMiON is now limited to using only $2\times 5\times 5$ uniform quadrature points, because of the limited parameter budget of $3\cdot 10^{4}$ trainable parameters. The inductive bias presented in Section 5.1.2 is incorporated in the data NGOs, data-free NGOs and model NGOs. Details of the model architectures are highlighted in Appendix C.2. The models have been trained on dataset D, described in Appendix B.2, using the training procedure described in Appendix D.2.
@@
Figure 17: Single time step relative $L^{2}$ test error versus (a) time scale $\tau/T$ and (b) length scale $\lambda/L$ for a standalone NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO. All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO) use the same 2 linear $\times$ 10 cubic $\times$ 10 cubic B-spline basis in space-time $(t,x,y)$. Points and error bars are, respectively, averages and 95% confidence intervals on 1000 manufactured solutions generated in the same way as dataset D, but with $\lambda/L=1$ (in (a)) and with $\tau/T=1$ (in (b)) (where we used $L=T=1$).
@@
Figure 18: Relative time step $L^{2}$ test error over time on (a) in distribution data with time and length scale $\tau/T=\lambda/L=1$ and (b) out of distribution data with $\tau/T=\lambda/L=0.1$, for a standalone NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO (an FNO is also added for comparison). All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO) use the same 2 linear $\times$ 10 cubic $\times$ 10 cubic B-spline basis, and are thus limited by the same $L^{2}$ projection error lower bound, indicated in black. Points and shaded areas are, respectively, averages and 95% confidence intervals on 100 manufactured solutions generated in the same way as dataset D, but with $\tau/T=\lambda/L=1$ (in (a)) and with $\tau/T=\lambda/L=0.1$ (in (b)) (where we used $L=T=1$).